OpenStudy (anonymous):

In the savannah, a sinkhole opened up catching some ostrich and cheetah. There was a total of 13 animals to be rescued. If 44 individual feet were pulled from the muck, how many ostrich and how many cheetahs were rescued? Part 1: Set up the problem. (2 points) Part 2: Explain why you set it up that way. (1 point) Part 3: Provide the solution AT LEAST SET UP THE PROBLEM PLEASE?

OpenStudy (anonymous):

a=number of ostrich, b=number of cheetah a+b=13 ==> b=13-a 2a+4b=44 2a+4(13-a)=44 2a+52-4a=44 -2a=-8 a=4 4+b=13 b=9